General Meeting Summary 8/23/17

From cua_phy
Jump to navigation Jump to search

DETECTORS AT JLAB (Arthur, Richard)

  • SHMS Aerogel detector
  • broken bolts on top fixed with help of Walter's team - uses same holes as before, drilled/tapped and installed new hex bolt, anticipate this to be a recurring issue due to materials difference (stainless steel bolt and aluminum box)
  • still need to take cosmic data to check if light leak fixed - ideally would like scintillators and drift chambers on during this test
  • Trays SP-20, SP-15 and SP-11 are stored in ESB building in roped off area near HKS equipment
  • Neutral Particle Spectrometer
  • Diameter is less than 19.2 mm as requested, result from 4 randomly selected PMTs: 19.0 +-0.2 mm - see this picture
  • Labels (est. thickness 50-100 um) are not glued to PMT body as requested to maximize space for shielding - see picture
  • Crystal arrival - latest update: delivery of all 460 by end of September
  • Magnet main coil shipment delayed until end of November (A few pictures: Pic1, Pic2)
  • HMS Aerogel Detector
  • Testing method to remove Bates PMTs without breaking them
  • Working on PMT scanning setup to test Bates PMTs once installed in the cylinders


PBWO4 STUDIES AT CUA (Richard, Andres)

  • design drawings including thermal design complete
  • indicate existing holes on furnace sides
  • add additional holes needed on furnace top and sides
  • indicate furnace heating elements for interference checks
  • mechanical setup ready and calibrated - link the results
  • heating elements have been installed
  • thermal profile simulation - initial results available
  • Crystal testing
  • all systems ready for crystal testing
  • commissioning of new pulse generator ongoing
  • possible incompatibility between TTL output and electronics - check out manual for input pulse requirements
  • reflector optimization studies ongoing


DATA ANALYSES AND SIMULATIONS (Salina, Dannie)

  • need to resolve issue of current dependence of DVCS coda events vs. DIS events
  • investigate calo part (bottom circuit on page 2) vs. HRS singles (top part of same diagram)
  • Note that HRS has delay of ~1 usec to form trigger, while ARS (calo) takes about 600 ns for signal


  • Resolved several issues in uncertainty calculation, e.g. dependence of mean on specific binning and unit inconsistencies
  • Results look reasonable and encouraging for structure function studies!
  • may want to check consistency with Kijun's calculation of uncertainties
  • Work on detector simulations ongoing


OTHER

  • Vladimir still waiting for visa - hopefully after 1 Sept


NEXT MEETING: WED 30 AUGUST at 10:00 AM (EDT)